OpenStudy (anonymous):

find the inverse of the function: g(x)=2x^2+6x+2

OpenStudy (anonymous):

hahahaha

terenzreignz (terenzreignz):

hirap naman nito

OpenStudy (anonymous):

waah, grabeh, xDD andali lang nian ehh

terenzreignz (terenzreignz):

ganun ba? pa demo nga, master...

OpenStudy (anonymous):

awtzs, nakakahiya namans a college diyan diba?? xDDDD

terenzreignz (terenzreignz):

magaling ka naman eh :P lalabas muna ako... pagbalik ko dapat nai-post mo na yung sagot mo... Derek-sensei :D

OpenStudy (anonymous):

waah, xDD not japanese. xDDD

terenzreignz (terenzreignz):

Sifu-Derek

hartnn (hartnn):

i am so lost :P

terenzreignz (terenzreignz):

terenz: this is so hard derek: waaah 'grabeh' (grave?) this is easy (eh) terenz: is that so? demonstrate, master derek: awtzs, wouldn't want to put the college student to shame, no? terenz: you're so good after all? I'll go out for a while, when I go back, you better post your answer, derek-sensei That's the gist of it @hartnn :)

hartnn (hartnn):

lol, thank you!

OpenStudy (anonymous):

filipino -english translator , ahahah, xD

hartnn (hartnn):

back to solving this ?

hartnn (hartnn):

i guess completing the square will help....

OpenStudy (anonymous):

@mathslover

terenzreignz (terenzreignz):

completing the square.... you know how to do it?

OpenStudy (anonymous):

yup, (b/2)^2

terenzreignz (terenzreignz):

It would have been more polite to respond to @hartnn :D anyway, first divide both sides of \[\Large y = 2x^2 + 6x +2 \] by 2 what do you get?

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

y/2=2x^2+6x+2/2

terenzreignz (terenzreignz):

yeah... so? simplify remember \[\large 2x^2 + 6x + 2 = 2(x^2 + 3x + 1)\]

OpenStudy (anonymous):

yup,, xD

terenzreignz (terenzreignz):

Well? Simplify. \[\Large \frac{y}2 = \color{red}?\]

OpenStudy (anonymous):

y/2

OpenStudy (anonymous):

y/2=x^2+3x+1

OpenStudy (anonymous):

@terenzreignz ,, busy????

terenzreignz (terenzreignz):

Overreacting much? :P You're the one that took long to answer my question :P Anyway, you're right. Now complete the square on the right side... using whatever technique was taught to you.

OpenStudy (anonymous):

huh?? panu ma complete ung square if may c na sia

terenzreignz (terenzreignz):

<sarcastic tone> oo nga no? ilipat mo kaya yung 'c' sa kabila? >:) <end sarcastic tone>

OpenStudy (anonymous):

hahaah, xDDDD

terenzreignz (terenzreignz):

.. fine... -.- \[\Large \frac{y}2 - 1 = x^2+3\] o... masaya ka na? :D

OpenStudy (anonymous):

nagets ko naman ehh. -____-

terenzreignz (terenzreignz):

syempre hinde, may typo ako eh :D \[\Large \frac{y}2 - 1 = x^2+3\color{red}x\]

OpenStudy (anonymous):

y/2-1+x^2+3+9/4

terenzreignz (terenzreignz):

+? o =? y/2 -1 = ...blabla

terenzreignz (terenzreignz):

\[\Large \frac{y}2 - 1 \color{red}{+\frac94}= x^2 + 3\color{green}x \color{red}{+\frac94}\] Wag kalimuta na kailangan sa PAREHONG side i-add yung 9/4... para patas naman, diba? :P

terenzreignz (terenzreignz):

*kalimutan

OpenStudy (anonymous):

aiyt, ahahah, xD nag crash ung broswer, ou nga noh,??? xD ahahah di ko napansin

terenzreignz (terenzreignz):

o... pano simplify mo yung left-side, tapos i-express mo yung right-side bilang perfect square.

OpenStudy (anonymous):

wait lang, xDD diba kailangan ntin i change ung x and y???

OpenStudy (anonymous):

ung sa pinaka start??

terenzreignz (terenzreignz):

di i-change mo. kung saan ka masaya suportahan taka \[\Large \frac{x}2 - 1 \color{red}{+\frac94}= y^2 + 3\color{green}y \color{red}{+\frac94}\]

OpenStudy (anonymous):

ahahahah, xDDDD

terenzreignz (terenzreignz):

You must understand, switching x and y is just a formality. In fact, it wouldn't make a difference in essence if we keep them as is, and do the changing afterwards.

OpenStudy (anonymous):

okay ,,,

terenzreignz (terenzreignz):

so... simplify the left, express the right as a perfect square... let's go now... vamos!

OpenStudy (anonymous):

2x+5/4

terenzreignz (terenzreignz):

galing naman ni kuya :D yung kanan naman :P

OpenStudy (anonymous):

hala. xDDD

OpenStudy (anonymous):

di ako marunong sa right side. xDDDD

OpenStudy (anonymous):

(y+3/2)(y+3/2)??

terenzreignz (terenzreignz):

O, kaya mo naman pala eh ... -.- \[\Large x^2 +bx + \frac{b^2}4 = \left(x+\frac{b}2\right)^2\]

terenzreignz (terenzreignz):

\[\Large \frac{2x+5}{4}= \left(y+\frac32\right)\left(y+\frac32\right)=\left(y+\frac32\right)^2\]

OpenStudy (anonymous):

tapos???

terenzreignz (terenzreignz):

tápos? tapós! \[\Large \frac{2x+5}{4}= \left(y+\frac32\right)^2\] Don't be shy, get the square root of both sides...

OpenStudy (anonymous):